Please help me find the following:
csc60
sec(-2pi/3)
and
cot-135
Thank you

Answer by Alan3354(21607)

Please help me find the following:
csc60
csc = 1/sin
sin(60) = sqrt(3)/2
csc(60) = 2/sqrt(3) = 2sqrt(3)/3
------------------------
sec(-2pi/3)
sec = 1/cos
cos(-2pi/3) = cos(pi/3) = 1/2
sec = 2
--------------
and
cot-135
cot = 1/tan
tan(-135) = 1
cot = 1